OpenStudy (anonymous):

How do I factorise 12a^2 – 11a – 15 = 0

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. 12a2−11a−15=0 Step 1: Factor left side of equation. (4a+3)(3a−5)=0 Step 2: Set factors equal to 0. 4a+3=0 or 3a−5=0 a= −3 4 or a= 5 3 Answer: a= −3 4 or a= 5 3

OpenStudy (nethercreep333):

-3/4 or a is 5/3

OpenStudy (anonymous):

i don't understand how you got (4a+3)(3a−5)=0

OpenStudy (nethercreep333):

Let's solve your equation step-by-step. 12a2−11a−15=0 Step 1: Factor left side of equation. (4a+3)(3a−5)=0 Step 2: Set factors equal to 0. 4a+3=0 or 3a−5=0 i dont know how to explain it any other way

OpenStudy (nethercreep333):

Im sorry maybe try khan academy and watch a video

OpenStudy (anonymous):

oh i think i understand now!

OpenStudy (nethercreep333):

can u medal me

OpenStudy (nethercreep333):

thanks lemme know if u need more help just tag me @nethercreep333

OpenStudy (anonymous):

i do have another question actually

OpenStudy (nethercreep333):

ok post

OpenStudy (anonymous):

Factorise 5x^2 – 9x – 2